OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope 1/2 and y-intercept (0, -5) and graph the line. Please show all of your work.

OpenStudy (anonymous):

y=mx+c is your standard equation where: m=slope= 1/2 c = y-intercept = -5 Therefore the equation will be... ?

OpenStudy (anonymous):

y=1/2x-5